### v1

Abstract
Large language models (LLMs) have achieved remarkable success and demonstrated superior performance across various tasks, including natural language processing (NLP), weather forecasting, biological protein folding, text generation, and solving mathematical problems. However, many real-world data exhibit highly non-Euclidean latent hierarchical anatomy, such as protein networks, transportation networks, financial networks, brain networks, and linguistic structures or syntactic trees in natural languages. Effectively learning intrinsic semantic entailment and hierarchical relationships from these raw, unstructured input data using LLMs remains an underexplored area. Due to its effectiveness in modeling tree-like hierarchical structures, hyperbolic geometry – a non-Euclidean space – has rapidly gained popularity as an expressive latent representation space for complex data modeling across domains such as graphs, images, languages, and multi-modal data. Here, we provide a comprehensive and contextual exposition of recent advancements in LLMs that leverage hyperbolic geometry as a representation space to enhance semantic representation learning and multi-scale reasoning. Specifically, the paper presents a taxonomy of the principal techniques of Hyperbolic LLMs (HypLLMs) in terms of four main categories: (1) hyperbolic LLMs through exp/log maps; (2) hyperbolic fine-tuned models; (3) fully hyperbolic LLMs, and (4) hyperbolic state-space models. We also explore crucial potential applications and outline future research directions. A repository of key papers, models, datasets, and code implementations is available at  https://github.com/sarangp2402/Hyperbolic-LLM-Models/tree/main.
keywords:
Large language models (LLMs), hierarchical structure, non-Euclidean manifold, hyperbolic geometry, semantic entailment, multi-resolution reasoning
1 Introduction
The emergence of foundation models and large language models (LLMs) has significantly empowered modern artificial intelligence (AI) systems, producing substantial impacts across a wide range of application fields, such as healthcare [52, 1], finance [64], computer programming [14, 110], earth system modeling [7, 99], autonomous systems [113], etc. Prevalent LLMs, such as GPT [2], Llama [33], Gemini [98], Claude [4], BERT [53], built on transformer architectures [105], excel at capturing long-range dependencies and perform effectively in various natural language processing (NLP) tasks, such as summarization, answering questions, translation, and generation of coherent and contextually relevant text. Typically, LLMs were designed to learn token embeddings in flat Euclidean space $\mathbb{R}^{n}$, where each token is represented as a single vector, leveraging the attention mechanism to capture rich contextual information. However, Euclidean embeddings present two major limitations: 1) flat Euclidean space (with curvature $\mathcal{K}=0$) has limited capacity to embed large and complex taxonomies with high distortions; 2) they struggle to capture logical and semantic entailment information during representation learning, including latent hierarchical structural features, which are critical for complex semantic reasoning and question answering tasks. This limitation is particularly evident when processing complex multi-modal data that exhibits tree-like structures with multi-scale features at various levels. Furthermore, numerous real-world datasets inherently exhibit hierarchical tree-like structures in non-Euclidean space [77, 57], such as textual entailment ontology data, linguistic structures, and phylogenetic trees of DNA sequences [8]. However, hyperbolic space ($\mathbb{H}^{n}$) with negative curvature ($\mathcal{K}<0$) is well-suited for representing the aforementioned large-scale hierarchical datasets, offering significant advantages over conventional Euclidean space ($\mathbb{R}^{n}$) for three key properties: 1) Hyperbolic space grows exponentially, aligning naturally with the tree-like expansion of hierarchical data, allowing for the effective preservation and representation of hierarchies with lower distortion. 2) Hyperbolic embeddings excel at preserving both local (leaf-level) and global (root-level) relationships, ensuring a comprehensive representation of hierarchical patterns. 3) Hyperbolic space ($\mathbb{H}^{n}$) offers high representational capacity even in low dimensions, reducing computational overhead.
As such, learning high-quality feature representations in hyperbolic space has gained significant interest across various research fields, including graph machine learning, computer vision, and natural language processing. Notably, in the domain of geometric deep learning (GDL), most complex networks are scale-free, characterized by a power-law distribution of node degrees. Hence, prevalent graph neural network (GNN) architectures have been successfully generalized to their hyperbolic variants, including hyperbolic node embedding with shallow skipgram model [11], hyperbolic GNNs [69], hyperbolic graph convolutional networks (GCNs) [13], and fully hyperbolic GNNs [108]. Hyperbolic GNN variants provide a powerful framework for modeling node hierarchies and embedding large-scale graphs into hyperbolic spaces. These approaches, characterized by fewer parameters and enhanced generalization capabilities, have been increasingly applied to various graph representation learning tasks, achieving great performance improvements over traditional Euclidean embeddings. Key applications include knowledge graph completion [57, 55, 12, 109, 114], and brain network embedding for detecting neurological degenerative diseases and predicting the age of the human brain [5, 90]. Yang et al. [121] provide a comprehensive review of hyperbolic graph neural network methods and their applications across multiple domains. However, despite these advances in hyperbolic embeddings for specific domains, a comprehensive understanding of how hyperbolic geometry can be systematically integrated into large language models remains fragmented across the literature.
In the NLP domain, language data typically exhibits a Zipfian distribution, as described by Zipf’s law [85], which reflects its hierarchical structure and motivates the use of hyperbolic geometry for representation. Petrovski [84] was the first to demonstrate that embedding sentences via simple Möbius averaging in the Poincaré disk outperforms Euclidean averaging on standard textual-entailment benchmarks, showing that negative curvature can benefit natural-language inference even with shallow RNN encoders. To capture useful word-level hierarchies for a variety of NLP tasks, hyperbolic geometry (i.e., Poincaré ball model) has been used to enhance conventional word embedding methods (e.g., word2vec and GloVe) [77, 100]. In particular, hyperbolic word embeddings have demonstrated superior performance compared to their Euclidean counterparts in lexical entailment prediction, achieving greater efficiency and improved generalization performance. Furthermore, hyperbolic word embeddings have demonstrated exceptional performance in word similarity, analogy, and hypernymy detection by using Gaussian mappings [100] in the Riemannian manifold. To extend hyperbolic embedding approaches to “deep” sequence models such as recurrent neural networks (RNNs), [39] first proposed generalizing vanilla RNNs and gated recurrent units (GRUs) to hyperbolic spaces, where the authors leveraged the theory of gyrovector spaces to reformulate Euclidean softmax regression, feedforward layers, RNNs, and GRUs, enabling them to operate on the Poincaré ball. The results demonstrate significant improvements in the performance of natural language inference (e.g., textual entailment) and noisy-prefix recognition. Fully hyperbolic neural networks [16] for NLP tasks represent a significant advancement by operating entirely within hyperbolic space using Lorentz transformations, overcoming the limitations of previous hybrid approaches that relied on tangent space operations.
Moreover, recent empirical investigations have revealed that pre-trained language models exhibit inherent hyperbolic characteristics that strongly support geometric approaches to language representation. For example, token frequency distributions in LLMs follow power-law patterns with exponents around $\alpha\approx 1.9$, where high-frequency abstract tokens cluster near the origin while low-frequency specific terms are positioned farther away, and the neural networks-derived token embeddings demonstrate significant hyperbolicity with hyperbolicity values of $\delta\approx 0.08-0.12$ across multiple LLM architectures [119]. These findings provide compelling empirical evidence that hyperbolic geometry aligns naturally with the intrinsic structure of language model representations, validating the theoretical foundations for hyperbolic LLMs development.
Herein, we primarily focus on recent advances in developing powerful LLMs that leverage hyperbolic geometry as their representation space, aiming to further enhance expressiveness, improve reasoning, capture hierarchical structures, and optimize efficiency for large-scale complex data. While the dominant paradigm has been Transformer-based architectures like GPT-4, Llama-3, Gemini, Claude-3, and BERT [105], which excel as LLMs, the field has also witnessed the emergence of alternative architectures such as state-space models (SSMs), with Mamba being a prominent example that offers improved computational efficiency over long-range dependencies [41]. Mamba’s improved successor Mamba2 [26] has also shown strong performance implementing state-space duality between attention and state space modeling. Both Transformer and Mamba architectures present unique opportunities and challenges when adapted to hyperbolic geometry, which we explore throughout this paper. We use the term LLMs broadly to encompass modern foundation models across diverse architectures and modalities, not limited to language data, but on extensive datasets to achieve strong performance across multiple downstream tasks. In Section 2, we first present the foundations of hyperbolic geometry, including hyperbolicity measures, hyperbolic models, and key optimization techniques in the hyperbolic manifold. Section 3 presents a taxonomy of existing hyperbolic LLMs approaches. Section 4 provides a benchmarking analysis of hyperbolic LLMs, evaluating their performance on mathematical reasoning, hierarchical language reasoning, while identifying current benchmarking limitations and the need for standardized evaluation frameworks. Section 5 explores emerging applications of hyperbolic LLMs across three primary domains: natural language processing, computer vision, and multi-modal visual-semantic representation and inference. Finally, we conclude the paper with a discussion on future research directions.
2 Foundations of hyperbolic geometry
Hyperbolic space $\mathbb{H}^{n}$ is a non-Euclidean geometric space with constant sectional negative curvature ($\mathcal{K}<0$), in which the parallel postulate fails, i.e., allowing multiple parallel lines to pass through a point outside a given line. Hyperbolic space can be seen as a “continuous version of trees”, characterized by negative curvature that makes distances grow exponentially toward the boundary. Notably, hyperbolic space shares a deep connection with Minkowski spacetime in special relativity, as the Lorentz model of hyperbolic geometry is isometric to the hyperboloid model used in Minkowski space [6, 78]. Figure 1 illustrates the three fundamental geometric spaces that underpin different representation learning approaches. Euclidean space ($\mathcal{K}=0$) serves as the foundation for traditional LLMs due to its computational simplicity and compatibility with standard linear algebra operations. Spherical geometry ($\mathcal{K}>0$) finds applications in constraining embeddings to unit spheres and modeling globally compact topologies. Hyperbolic space ($\mathcal{K}<0$), distinguished by its exponential volume growth, provides the geometric foundation for HypLLMs. Moreover, hyperbolic embeddings can represent hierarchical structures much more parsimoniously – for instance, Nickel and Kiela [77] showed that a 2-dimensional hyperbolic model captured the WordNet noun hierarchy better than a 200-dimensional Euclidean model, highlighting the vastly greater representational capacity of negatively curved space.
Euclidean ($\mathcal{K}=0$) Can be used for generic data Spherical ($\mathcal{K}>0$) Can be used for globally compact, closed topology Hyperbolic ($\mathcal{K}<0$) Can be used for hierarchical, tree structures
Figure 1: Three commonly used representation spaces: Euclidean space, where traditional LLMs operate; spherical geometry, which constrains embeddings to bounded surfaces; and hyperbolic space, the basis of HypLLMs, which naturally models hierarchical structures.
2.1 Hyperbolicity measure
$\delta$-hyperbolicity introduced by Gromov [40] quantifies how closely a space resembles a tree-like structure, where $\delta\geq 0$ measures the deviation from perfect tree-like geometry. For any 4-tuple $(x,y,z,t)$ in a metric space, $\delta(x,y,z,t)$ is defined as half the difference between the biggest two of the following sums:
(1) $d(x,y)+d(z,t),\quad d(x,z)+d(y,t),\quad d(x,t)+d(y,z).$
The space’s $\delta$-hyperbolicity is defined as the supremum of these values across all possible 4-tuples. Smaller $\delta$ values indicate a higher similarity to tree structures, making this hyperbolic space more naturally suited for hierarchical modeling [77, 38].
2.2 Hyperbolic representation models
Two commonly used isomorphic models for representing hyperbolic space, the Poincaré ball ($\mathcal{B}_{r}^{n}$) model and the Lorentz model (or hyperboloid manifold - $\mathcal{L}_{r}^{n+1}$), are illustrated in the Figure 2. The sectional curvature of hyperbolic space is determined by $\mathcal{K}=-1/r^{2}$, where $r>0$ is a scale parameter (or radius) that determines the amount of curvature, i.e., larger $r$ $\rightarrow$ smaller curvature ($\mathcal{K}$) $\rightarrow$ flatter space; $\mathcal{K}=0$ is Euclidean space when $r\rightarrow\infty$. In the standard hyperbolic space with sectional curvature $\mathcal{K}=-1$, the curvature parameter $r=1$. Both $\mathcal{B}_{r}^{n}$ and $\mathcal{L}_{r}^{n+1}$ represent the same abstract $n$-dimensional hyperbolic space $\mathbb{H}^{n}$ with radius $r$, but use different coordinate systems. The Poincaré ball $\mathcal{B}_{r}^{n}$ provides a bounded $n$-dimensional representation, while the Lorentz model $\mathcal{L}_{r}^{n+1}$ embeds the hyperbolic space as a hyperboloid in $(n+1)$-dimensional Minkowski spacetime.
Figure 2: Illustration of the two primary hyperbolic representation models used in HypLLMs: the Lorentz model (top left) and the Poincaré ball model (right). The Lorentz model represents hyperbolic space on a single sheet of hyperboloid embedded in Minkowski space $\mathcal{L}_{r}^{n+1}$, offering closed-form geodesic computations and improved numerical stability. The Poincaré ball in space $\mathcal{B}_{r}^{n}$ is a conformal model where geodesics are circular arcs and points near the boundary encode fine-grained leaf-level hierarchy. The lower oval shows the tangent space used in log/exp-based manifold operations for optimization and model computation. These geometric representations provide the foundation for hierarchical embeddings in large-scale hyperbolic LLMs.
(1) Poincaré ball model
The Poincaré ball model represents $\mathbb{H}^{n}$ as the open unit ball $\mathcal{B}_{r}^{n}=\{x\in\mathbb{R}^{n}:\|x\|<r\}$ (where $\|\cdot\|_{2}$ denotes the Euclidean norm) with metric $ds^{2}=\lambda_{x}^{2}\,d\mathbf{x}^{2}$, where $d\mathbf{x}^{2}$ is the Euclidean metric and $\lambda_{x}=\frac{2}{1-\|x\|^{2}}$ is the conformal factor that scales distances according to the hyperbolic geometry [77]. Geodesics in this manifold are arcs that tend toward the boundary $\|x\|\to 1$. The geodesic distance between two points $u,v\in\mathbb{B}^{n}$ is:
(2) $d_{\mathbb{B}}(u,v)\;=\;\mathrm{arcosh}\!\Big{(}1+2\,\frac{\|u-v\|^{2}}{(1-\|u\|^{2})(1-\|v\|^{2})}\Big{)},$
which grows without bound as $u$ or $v$ approach the boundary of the ball. This model is advantageous for gradient-based optimization.
(2) Lorentz model
Alternatively, the Lorentz model represents $\mathbb{H}^{n}$ as a hyperboloid embedded in $\mathbb{R}^{n+1}$: $\mathcal{L}_{r}^{n+1}=\{x\in\mathbb{R}^{n+1}:\langle x,x\rangle_{L}=-r^{2},\;x_{0}>0\}$, equipped with the Minkowski inner product [78]:
(3) $\langle x,y\rangle_{L}=-x_{0}y_{0}+\sum_{i=1}^{n}x_{i}y_{i}.$
Note that the Lorentz model uses only the upper sheet ($x_{0}>0$) of the two-sheeted hyperboloid. This upper sheet constraint ensures that all points represent valid hyperbolic space coordinates. The hyperbolic distance in this model is
(4) $d_{\mathcal{L}}(x,y)\;=\;\mathrm{arcosh}\!\big{(}-\langle x,y\rangle_{L}\big{)},$
and one can convert between the Lorentz and Poincaré representations via explicit coordinate mappings [76]. Both models are useful; notably, the Lorentz model often provides better numerical stability for optimization (since it spreads points more evenly when far from the origin). In either model, gradient-based learning is performed with Riemannian optimization techniques (e.g., RSGD) or projected Euclidean updates via the exponential and logarithmic maps.
2.3 Exponential and logarithmic mapping
Ganea et al. [38] formalized hybrid operations for hyperbolic neural networks using logarithmic and exponential mappings, establishing a principled framework for bridging Euclidean and hyperbolic computations. This approach enables the adaptation of standard neural network operations to hyperbolic manifolds through a systematic mapping strategy. Specifically, it leverages the tangent space at the origin of the Poincaré ball, which is isometric to Euclidean space $\mathbb{R}^{n}$. The logarithmic map projects points from the hyperbolic manifold to this tangent space by
(5) $\text{log}_{0}(\mathbf{x})=\text{artanh}(\|\mathbf{x}\|)\frac{\mathbf{x}}{\|\mathbf{x}\|},$
with the inverse exponential map returning points to the hyperbolic space using
(6) $\text{exp}_{0}(\mathbf{v})=\tanh(\|\mathbf{v}\|)\frac{\mathbf{v}}{\|\mathbf{v}\|}.$
These mappings preserve distances in the radial direction and enable a workflow in which (i) hyperbolic points are mapped to the tangent space via $\text{log}_{0}$, (ii) standard Euclidean operations are performed, and (iii) results are mapped back to hyperbolic space via $\text{exp}_{0}$. To enable direct computation in hyperbolic space, Ungar [104, 103] introduced Möbius gyrovector algebra, which Ganea et al. [39] highlighted and utilized to generalize vector space operations to the Poincaré ball. The key operations include:
Möbius addition: For $\mathbf{x},\mathbf{y}\in\mathcal{B}_{r}^{n}$,
(7) $\mathbf{x}\oplus\mathbf{y}=\frac{(1+2\langle\mathbf{x},\mathbf{y}\rangle+\|\mathbf{y}\|^{2})\mathbf{x}+(1-\|\mathbf{x}\|^{2})\mathbf{y}}{1+2\langle\mathbf{x},\mathbf{y}\rangle+\|\mathbf{x}\|^{2}\|\mathbf{y}\|^{2}}.$
Möbius scalar multiplication:
(8) $r\otimes\mathbf{x}=\tanh(r\cdot\text{artanh}(\|\mathbf{x}\|))\frac{\mathbf{x}}{\|\mathbf{x}\|}.$
Möbius matrix-vector multiplication: For matrix $\mathbf{M}$ and hyperbolic point $\mathbf{x}$,
(9) $\mathbf{M}\otimes\mathbf{x}=\text{exp}_{0}(\mathbf{M}\cdot\text{log}_{0}(\mathbf{x})).$
These operations preserve the hyperbolic structure while enabling neural network computations such as linear transformations and activations.
2.4 Optimization techniques in Riemannian manifold
Training large models in hyperbolic space requires specialized optimization techniques, since standard Euclidean gradient descent does not directly apply to curved manifolds. Notably, numerical instabilities such as overflow, underflow, and NaNs can arise in both the Poincaré ball and Lorentz models when points approach the boundary or stray far from the origin. To combat these issues, researchers have proposed reparameterization tricks – for example, performing updates in Euclidean parameter space and mapping them into hyperbolic space – to keep representations numerically well-behaved. Furthermore, optimization algorithms have been extended to the hyperbolic realm: Riemannian SGD (RSGD) and Riemannian Adam (RADAM) update weights in the tangent space at each point on the manifold and then use the exponential map to retract back to hyperbolic space [38]. Ganea et al.[39] introduced the Möbius gyrovector framework (e.g., Möbius addition and scalar multiplication) to define neural network operations entirely in the Poincaré ball model, enabling the construction and training of hyperbolic feed-forward and recurrent networks. A key insight is to perform gradient updates intrinsically on the manifold via Riemannian optimization [68].
Riemannian Stochastic Gradient Descent (RSGD): Riemannian Stochastic Gradient Descent (RSGD) generalizes SGD to Riemannian manifolds by using exponential mappings to update parameters while maintaining manifold constraints [9]:
(10) $\theta_{t+1}=\exp_{\theta_{t}}(-\eta_{t}\nabla_{R}\mathcal{L}(\theta_{t})),$
where $\exp_{\theta_{t}}$ is the exponential map at parameter $\theta_{t}$ (point on the manifold), $\eta_{t}$ is the learning rate, and $\nabla_{R}\mathcal{L}(\theta_{t})$ is the Riemannian gradient of the loss function. Bonnabel first proved that RSGD updates converge under certain conditions on geodesically convex functions, establishing the theoretical foundation for hyperbolic optimization [9].
Riemannian Stochastic Variance Reduced Gradient (RSVRG): Building on this foundation, Zhang et al. [127] introduced Riemannian SVRG (RSVRG), the first variance-reduced stochastic optimization method for Riemannian manifolds. It applies the “Stochastic Variance Reduced Gradient (SVRG)” technique in non-Euclidean settings to accelerate convergence. RSVRG combines geometric insights with variance reduction techniques, achieving linear convergence for geodesically convex functions and $\mathcal{O}(1/T)$ convergence for non-convex objectives [127, 68].
Riemannian Adagrad (RADAGRAD): Recently, RADAGRAD [43] extended adaptive gradient methods to Riemannian manifolds. For a product manifold $\mathcal{M}=\mathcal{M}_{1}\times\cdots\times\mathcal{M}_{n}$, the parameter update becomes:
(11) $\theta_{t+1}^{(i)}=\exp_{\theta_{t}^{(i)}}\left(-\eta_{t}\frac{\nabla_{R}^{(i)}\mathcal{L}(\theta_{t})}{\sqrt{v_{t}^{(i)}+\epsilon}}\right),$
where $v_{t}^{(i)}$ tracks historical gradients for component $\mathcal{M}_{i}$. This adaptive approach demonstrates superior convergence properties compared to RSGD [43]. These techniques significantly reduce computational overhead while maintaining the benefits of hyperbolic geometry.
In practice, modern hyperbolic LLM frameworks leverage Riemannian Adam or Riemannian Adagrad [43] (extensions of popular Euclidean optimizers adapted to hyperbolic manifolds) to achieve stable training. The choice between the Poincaré ball and Lorentz models significantly impacts optimization stability. The Lorentz model enables more stable optimization through efficient closed-form geodesic computations, avoiding numerical instabilities from Poincaré distance calculations while preserving the exponential growth properties essential for hierarchical representations. Critical implementation details include gradient rescaling to prevent boundary violations and appropriate initialization strategies to maintain numerical stability throughout training [76, 125]. Empirically, Nickel and Kiela [77, 78] noted that using the manifold’s geometry during optimization yields higher-quality embeddings and faster convergence than naive Euclidean training with projection. Efficient training of hyperbolic neural networks remains an active area of research, with recent advances such as Sparse Spectral Training (SST) [130] offering scalable and memory-efficient optimization strategies for both Euclidean and hyperbolic models, including applications in node classification and link prediction.
2.5 Main challenges in building hyperbolic LLMs
Despite the theoretic-al advantages of hyperbolic geometry for modeling hierarchical data, developing practical hyperbolic large language models faces several fundamental challenges that significantly impact their scalability.
(1) Numerical stability and representation capacity
One of the most critical challenges in hyperbolic LLMs stems from the inherent numerical instabilities of hyperbolic representations in floating-point arithmetic. As demonstrated by Mishne et al. [76], the exponential volume growth property of hyperbolic space leads to catastrophic numerical problems in floating-point arithmetic.
Poincaré ball limitations: Under 64-bit floating-point arithmetic, the Poincaré model exhibits severe representation constraints. Points can only be accurately represented within a ball of radius $r_{0}\approx 38$ (where $r_{0}$ is the maximum stable radius under 64-bit floating-point precision), as the boundary proximity (when $\|\mathbf{x}\|\to 1$) causes numerical instability in the conformal factor $\lambda_{x}=\frac{2}{1-\|\mathbf{x}\|^{2}}$. This leads to gradient vanishing problems where the Riemannian gradient magnitude scales as $\mathcal{O}(\delta^{2})$ for points at distance $\delta$ from the boundary. As demonstrated by Mishne et al. [76], the maximum representable distance corresponds to $|\mathcal{K}|\approx 16$ in the equation $1-10^{-|\mathcal{K}|}=2^{-53}$, severely limiting the model’s capacity to represent distant hierarchical relationships.
Lorentz model constraints: The Lorentz model’s closed-form geodesic computation partially mitigates numerical issues by avoiding division by near-zero denominators in distance calculations:
(12) $d_{\mathcal{L}}(\mathbf{x},\mathbf{y})=\text{arccosh}(-\langle\mathbf{x},\mathbf{y}\rangle_{\mathcal{L}}),$
where $\langle\cdot,\cdot\rangle_{\mathcal{L}}$ denotes the Minkowski inner product. However, the Lorentz model still faces representation capacity limitations within the radius $r_{0}/2\approx 19$.
Sala et al. [92] rigorously characterize a fundamental precision–depth trade‐off for hyperbolic embeddings. They prove that embedding a simple chain graph of length $\ell$ with distortion at most $1+\varepsilon$ requires either: (i) scaling the effective hyperbolic radius linearly in $\ell$; or (ii) allocating $\Theta(\ell/\varepsilon)$ bits of floating-point mantissa precision. Thus, the need to balance numerical precision against the depth (chain length) of the hierarchy is an intrinsic property of hyperbolic geometry.
(2) Computational overhead and scalability
Hyperbolic LLMs face significant computational challenges that limit their practical deployment, particularly for large-scale applications.
Exponential and logarithmic map overhead: The frequent use of exponential and logarithmic mappings between hyperbolic space and tangent space introduces substantial computational overhead [15, 119]. Each transformation requires expensive operations including arcosh, sinh, and cosh functions, leading to training times that are slower than standard Euclidean models. For hybrid architectures, this overhead is amplified by the need for repeated mappings in attention mechanisms, feedforward layers, and normalization operations between Euclidean tangent space and hyperbolic embeddings.
Quadratic complexity in attention: Traditional hyperbolic attention mechanisms suffer from quadratic time complexity $\mathcal{O}(N^{2})$ with respect to the number of input tokens. This scalability bottleneck becomes particularly problematic for long-sequence modeling and large-scale graph data processing.
Memory and precision requirements: Hyperbolic operations often require higher precision arithmetic to maintain numerical stability, leading to increased memory consumption and reduced computational efficiency on hardware accelerators. Multi-component floating-point representations have been proposed as solutions [125], but these approaches introduce additional implementation complexity and may not be efficiently supported on all hardware platforms. To be precise, BigFloats required for computations proposed by Yu et al. [125] are not supported on ML accelerator hardware like GPUs: computations using BigFloats on CPUs will be slow and hence not realistic for many meaningful tasks. Secondly, BigFloats are currently not supported in most deep learning frameworks, even on CPUs.
(3) Training and optimization challenges
Training hyperbolic LLMs presents unique training and optimization challenges that require careful tuning of operations in hyperbolic spaces compared to the training operations happening in standard Euclidean models.
Limited hyperbolic operations: Maintaining points on the hyperbolic manifold throughout training is a critical challenge. Standard optimization steps can easily violate manifold constraints, requiring projection operations or constraint-preserving update schemes. These projections introduce additional computational overhead and can accumulate errors over many training iterations, particularly in deep networks with multiple hyperbolic layers. Many standard neural network operations lack well-defined hyperbolic counterparts. Operations such as batch normalization, dropout, concatenation, and various activation functions require careful geometric reformulation to preserve hyperbolic properties [120, 16]. The absence of these operations has historically limited hyperbolic models to simple architectures, hindering their application to complex tasks that benefit from deep, sophisticated networks.
Curvature selection and adaptation: Choosing appropriate curvature values for different layers and components remains a significant challenge. Fixed curvature models may not optimally represent data with varying hierarchical depths, while learnable curvature approaches introduce additional hyperparameters and optimization complexity [45]. Recent mixture-of-curvature approaches show promise but add substantial architectural complexity.
These challenges collectively represent significant barriers to the widespread adoption of hyperbolic LLMs. However, ongoing research continues to address these limitations through improved numerical methods, more efficient architectures, and better optimization techniques. Understanding these challenges is crucial for researchers seeking to develop robust and scalable hyperbolic language models that can realize the theoretical benefits of hyperbolic geometry in practical applications. Building upon these mathematical foundations, we now present a comprehensive taxonomy of hyperbolic LLM architectures and examine how these geometric principles have been integrated into large language model systems.
3 Hyperbolic LLMs
A major challenge in developing large language models (LLMs) in hyperbolic space lies in generalizing Euclidean operations to hyperbolic manifolds. Unlike simple element-wise operations (e.g., addition, multiplication, etc.) or fundamental neural network components such as linear transformations, attention mechanisms, and normalization, these operations and components require careful geometric reformulation in hyperbolic space. Even operations like convolution and pooling must be redesigned to respect the curved geometry of hyperbolic space, as standard Euclidean formulations do not preserve the geometric properties essential for hierarchical representation. The fundamental issue stems from the fact that hyperbolic manifolds do not naturally support the linear algebraic operations that underpin modern neural network architectures. This necessitates the development of specialized geometric operations and optimization techniques that respect the curved geometry of hyperbolic space while maintaining computational efficiency. Table 1 presents our taxonomy of these models across four primary categories: hybrid hyperbolic-Euclidean models, hyperbolic fine-tuned models, fully hyperbolic models, and hyperbolic state-space models. Table 2 provides a comprehensive overview of the tasks, dataset types, and specific datasets used to evaluate these HypLLM models across different application domains.
Table 1: Taxonomy of hyperbolic large language models (HypLLMs), organized by architecture type, backbone, model, and geometry.
Category Backbone Model Geometry
Hybrid hyperbolic- Euclidean models Transformer Hyperbolic BERT [15] Lorentz
Transformer HiT [47] Poincaré ball
Transformer PoinCLIP [94] Poincaré ball
Transformer Hyperbolic BLIP-2 [72] Poincaré ball
Transformer HyperLLM [18] Poincaré ball and Lorentz
Transformer HERec [70] Lorentz
Tri-Modal MoE HyperSurv [116] Lorentz
Attention-based NN HySurvPred [117] Poincaré ball
Attention-based NN ANTHEM [21] Lorentz
Hyperbolic fine-tuned models Transformer HypLoRA [119] Lorentz
Transformer HoRA [118] Lorentz
Fully hyperbolic models Transformer Hypformer [120] Lorentz
Transformer HELM [45] Lorentz
Transformer L-CLIP [46] Lorentz
GNN+Transformer HypGraphRAG [46] Lorentz
Hyperbolic state-space models Mamba SHMamba [123] Poincaré ball
Mamba HMamba-Full [128] Lorentz
Mamba HMamba-Half [128] Half Euclidean - Half Lorentz
Mamba-2 HiM-Poincaré [83] Poincaré ball
Mamba-2 HiM-Lorentz [83] Lorentz
Table 2: Representative hyperbolic large language models (HypLLMs) with their learning tasks, dataset types, and benchmark datasets.
Model Task Dataset type Datasets
Hyperbolic BERT [15] Entailment detection Language GAP [111], DPR [89], WSC [61], WG [91], PDP [27], FIGER [67], Open Entity [20], CoNLL-2003 [101], Few-NERD [31], TACRED [129], TACREV [3], Re-TACRED [96], MRQA [35]
HiT [47] Mixed-hop prediction, multi-hop inference Language WordNet [75], FoodOn [32], DOID [93], SNOMED [95]
PoinCLIP [94] Zero-shot image classification and retrieval Image-text pairs Food-101 [10], CIFAR-10 [58], CIFAR-100 [58], CUB-200-2011 [107], SUN397 [115], Aircraft [71], DTD [22], Pets [81], Caltech-101 [34], Flowers [79], STL-10 [24], EuroSAT [48], RESISC45 [17], Country211 [88], MNIST [59], CLEVR [51], PCam [106], SST2 [88]
Hyperbolic BLIP-2 [72] Image captioning, image-text retrieval Image-text pairs MS COCO [65]
HyperLLM [18] Recommendation User-item tabular data Amazon-Toys [73], Amazon-Sports [73], Amazon-Beauty [73]
HERec [70] Recommendation User-item graph data Amazon-Books [70], Yelp [70], Google-Reviews [70]
HyperSurv [116] Survival prediction Histological images-text pairs TCGA [102]
HySurvPred [117] Survival prediction Histological images-genomics pairs TCGA [102]
ANTHEM [21] Product search ranking E-commerce search E-commerce search data
HypLoRA [119] Arithmetic reasoning Math problems MAWPS [56], SVAMP [82], GSM8K [25], AQuA [66]
HoRA [118] Mathematical reasoning Math problems MAWPS [56], SVAMP [82], GSM8K [25], AQuA [66]
Hypformer [120] Graph node classification Graph Amazon2M [19], ogbn-proteins [50], ogbn-arxiv [50], ogbn-papers100M [50]
HELM [45] Multiple-choice QA QA benchmark Wikipedia [112], MMLU [49], ARC-C [23], CommonsenseQA [97], HellaSwag [126], OpenBookQA [74]
L-CLIP [46] Image-text retrieval Image-text pairs CIFAR-10 [58], CIFAR-100 [58], ImageNet [28], RedCaps [29]
HypGraphRAG [46] Knowledge QA Knowledge graphs WebQSP [124]
HiM-Poincaré [83] Mixed-hop prediction, multi-hop inference Ontology WordNet [75], FoodOn [32], DOID [93], SNOMED [95]
HiM-Lorentz [83] Mixed-hop prediction, multi-hop inference Ontology WordNet [75], FoodOn [32], DOID [93], SNOMED [95]
SHMamba [123] Audio-Visual QA Audio-video pairs MUSIC-AVQA [62], AVQA [122]
HMamba-Full [128] Sequential recommendation Sequential interaction data ML-1M [44], Texas, California, New York
HMamba-Half [128] Sequential recommendation Sequential interaction data ML-1M [44], Texas, California, New York
Now we address each category with the fundamental challenges of either adapting Euclidean neural operations to curved manifolds or direct embedding in the hyperbolic manifold through distinct strategies such as hybrid models maintaining compatibility through strategic mapping, fine-tuning methods adapting existing models with minimal architectural changes, or fully hyperbolic approaches redesigning all components for native curved space operation. Formally, given input embeddings $X\in\mathbb{R}^{B\times L\times d}$, where $B$ is the batch size, $L$ is the sequence length, and $d$ is the embedding dimension, the model outputs embeddings $Y\in\mathbb{R}^{B\times L\times d}$. We denote the hyperbolic embedding space as $\mathbb{H}^{n}$ with Euclidean space as $\mathbb{R}^{n}$. These notations will be used throughout Figure 3-5 to describe input/output flows and embedding spaces.
3.1 Hybrid hyperbolic-Euclidean models
As illustrated in Figure 3, hybrid hyperbolic-Euclidean models represent the most common approach to integrating hyperbolic geometry into large language models. These architectures maintain standard Euclidean operations in most components while strategically incorporating hyperbolic representations in specific layers or modules through exponential/logarithmic mappings. The typical workflow involves: (1) mapping Euclidean input embeddings $X\in\mathbb{R}^{B\times L\times d}$ to hyperbolic space $\mathbb{H}^{n}$ via exponential maps (Poincaré ball or Lorentz lifting), (2) performing geometric operations in the curved space, and (3) mapping back to Euclidean space through logarithmic maps for compatibility with standard transformer components.
Figure 3: Hybrid models use exponential/logarithmic mappings between Euclidean and hyperbolic spaces with options for Poincaré ball or Lorentz embeddings.
Researchers have begun to incorporate hyperbolic geometry directly into large language models to better capture latent hierarchies in language data. Chen et al. introduced Hyperbolic BERT [15], which represents a significant advancement in integrating hyperbolic geometry with pre-trained language models. The authors situate BERT entirely within hyperbolic space by mapping dependency trees with geodesic distances on the Lorentz model. While Hyperbolic BERT demonstrates superior performance on tasks requiring hierarchical reasoning, it operates approximately 1.3 times slower than standard BERT due to the added computational overhead of manifold operations; additionally, it can exhibit numerical instabilities during repeated transformations. The model processes input sequences by first embedding tokens in $\mathbb{R}^{512}$, then mapping to the Lorentz manifold $\mathbb{L}^{512}$ where attention computations leverage hyperbolic distances. Each transformer layer maintains this hybrid hyperbolic-Euclidean structure, with feed-forward networks operating in tangent spaces to preserve computational efficiency while capturing hierarchical relationships through curved geometry. This computational overhead highlights the delicate balance between geometric expressiveness and practical deployability in hyperbolic LLMs. Another notable effort is HiT [47], which extends Transformer-based LLMs to encode linguistic hierarchies in a Poincaré ball. HiT similarly embeds tokens in a hyperbolic latent space to capture parse-tree structures, reporting improved entailment and transitive reasoning. However, like Hyperbolic BERT, this approach incurs a non-trivial slowdown on long sequences due to the costly curvature-aware operations. In a multimodal context, a related approach, PoinCLIP [94] projects CLIP’s image-text embeddings into a Poincaré ball to enforce hierarchical distances, yielding improved zero-shot classification on coarse-grained categories.
In summary, the exp/log map techniques in these models integrate hyperbolic geometry within the model’s forward pass (at multiple layers), rather than only at the input or output embedding layer, effectively making the network partially hyperbolic. Having established the mathematical foundations for hybrid hyperbolic-Euclidean models, we now also examine parameter-efficient approaches that adapt pre-trained LLMs to hyperbolic space through targeted fine-tuning strategies.
3.2 Hyperbolic fine-tuned models
Recent research has explored methods for fine-tuning pre-trained LLMs in hyperbolic space to better capture hierarchical structures without retraining entire models from scratch. Hyperbolic fine-tuned models represent a special type of hybrid hyperbolic-Euclidean models (see section. 3.1) that adapt frozen pre-trained architectures to hyperbolic space rather than integrating hyperbolic operations within the forward pass. As depicted in Figure 4, hyperbolic fine-tuned models represent a parameter-efficient approach to adapting pre-trained LLMs for hyperbolic geometry. The architecture centers around a frozen LLM backbone with specialized hyperbolic adapters that perform curvature-constrained updates. This approach leverages the observation that pre-trained token embeddings often exhibit inherent hierarchical structure, making them suitable candidates for fine-tuning a Euclidean model in a hyperbolic space without full model retraining.
Figure 4: Fine-tuned approaches adapt frozen pre-trained models with hyperbolic adapters using curvature-constrained updates and gradient clipping.
(1) HypLoRA: Yang et al. [119] observe that token embeddings from a pre-trained GPT model exhibit a high degree of inherent hyperbolicity with frequent tokens clustering near the origin and rare tokens farther out, indicating a latent tree structure. To exploit this, they propose HypLoRA as a parameter-efficient fine-tuning method that avoids cancellation effects from naive exponential/logarithmic map applications. While claiming to operate on the Lorentz manifold, HypLoRA implements adapter-like modifications through performing tangent space projections and applying curvature-constrained weight perturbations in a Euclidean subspace.
Formally, given a hyperbolic embedding $x_{H}\in\mathbb{L}^{n}$ and a weight matrix $W$, one maps to the Euclidean tangent space via $log_{0}(x_{H})$, applies the linear transformation, and then maps back using the exponential map: $\exp_{0}\big{(}W\log_{0}(x_{H})\big{)}.$ Incorporating a LoRA update $\Delta W=BA$, the transformation becomes: $\exp_{0}\big{(}(W+BA)\log_{0}(x_{H})\big{)},$ which, after inverse mapping and approximation, corresponds to the Euclidean expression: $Wx_{E}+BAx_{E}.$ HypLoRA avoids this by introducing a hyperbolic adapter: it maps $x_{E}$ to $x_{H}$ once, applies a low-rank update $\Delta W_{H}$ directly on the manifold, and maps back to Euclidean space as $x_{E}$ for compatibility with downstream layers. Despite demonstrating improvement on the AQuA dataset, HypLoRA’s fundamental limitation is that the base model remains Euclidean, with hyperbolic operations confined to adapter modules.
(2) HoRA: Building on similar principles as HypLoRA, Yang et al. [118] developed HoRA using curvature-scaled transformations:
(13) $\Delta W=\text{exp}_{0}(\alpha\cdot\underbrace{\text{log}_{0}(W_{\text{base}})}_{\text{Euclidean}}),$
where $\alpha$ learns layer-specific adaptation rates. Here, $\log_{0}$ and $\exp_{0}$ denote mappings to and from the Lorentz manifold (origin at 0). Intuitively, HoRA multiplies the Euclidean base weight by a factor $\alpha$ within hyperbolic space and then projects it back. This curvature-aware scaling yields the adapted weight $W_{\text{HoRA}}=W_{\text{base}}+\Delta W$. This adapter module operate with learnable curvature parameters, allowing the model to adaptively determine the optimal level of hierarchy for different linguistic structures.
While HoRA achieves 17.30% improvement over Euclidean LoRA on mathematical reasoning tasks, HoRA’s hyperbolic component is essentially a wrapper around Euclidean weights. Notably, these hyperbolic fine-tuning methods have demonstrated significant gains on complex reasoning tasks (e.g., math word problem solving). The “hyperbolic vanishing gradient” problem (where gradients in hyperbolic space can attenuate due to the curvature) may persist due to residual Euclidean computations in backpropagation, as the curvature-scaled training provides only partial mitigation of gradient attenuation in hyperbolic space. The success of these approaches validates the hypothesis that pre-trained models possess a latent hierarchical structure that can be effectively exploited through hyperbolic fine-tuning. While fine-tuning approaches provide computational efficiency, they maintain fundamental Euclidean operations. To fully leverage hyperbolic geometry’s representational power, we next explore architectures that operate entirely within the hyperbolic space.
3.3 Fully hyperbolic models
Figure 5 illustrates the architecture of fully hyperbolic models, which represent the most theoretically complete approach to hyperbolic LLMs. Unlike hybrid models that rely on exponential/logarithmic mappings, these architectures operate entirely within hyperbolic space $\mathbb{H}^{n}$ using specialized geometric operations. The workflow shows how input embeddings $X\in\mathbb{H}^{B\times L\times d}$ are processed through fully hyperbolic transformer blocks containing linear-time hyperbolic attention, Möbius multi-head attention or cone attention mechanisms, and Lorentz linear transformations, ultimately producing hyperbolic outputs $Y\in\mathbb{H}^{B\times L\times d}$.
Figure 5: Fully hyperbolic models operate entirely in curved space with specialized geometric operations including linear-time hyperbolic attention and Lorentz linear transformations.
While early models applied hyperbolic geometry only to partial components-e.g., replacing attention scores with negative hyperbolic distance [42]-recent approaches have extended curvature to the entire network. Notably, Yang et al. [120] proposed Hypformer using specialized hyperbolic transformation blocks (HTC and HRC) that operate directly on the Lorentz manifold, avoiding frequent exponential/logarithmic mappings between tangent and hyperbolic spaces. Hypformer achieves linear complexity through hyperbolic factorization of the attention mechanism. The traditional quadratic attention $\mathcal{O}(N^{2})$ is reduced to linear $\mathcal{O}(N)$ by changing the computation order from $(Q^{T}K)V$ to $Q(K^{T}V)$, where $Q$ is the query, $K$ is the key, and $V$ is the value:
(14) $\text{Attention}(Q,K,V)\approx\phi(Q)\psi(K)^{T}V,$
where $\phi(\cdot)$ and $\psi(\cdot)$ are hyperbolic feature maps that preserve the Lorentzian structure while enabling efficient computation through kernel approximation methods. The model defines hyperbolic analogues of all key layers, including linear projections, normalization, and activation functions. This linear-time “hyperbolic attention” mechanism eliminates the computational bottleneck of traditional quadratic attention while preserving the geometric properties essential for hierarchical modeling. Hypformer achieves superior results over both Euclidean and hybrid hyperbolic-Euclidean transformers on hierarchical reasoning tasks.
He et al. introduce HELM [45], a Hyperbolic LLM that operates “fully in hyperbolic space” while employing a Mixture-of-Curvature Experts design. Instead of using a single constant curvature for the entire model, HELM assigns different transformer experts to different curvature spaces (including hyperbolic subspaces with varying radii), allowing the model to adaptively represent features at multiple scales of hierarchy. A key innovation in HELM is the hyperbolic multi-head latent attention (HMLA) mechanism, which enables efficient reduced-KV-cache training and inference while preserving hyperbolic properties. HELM’s Mixture-of-Curvature Experts (MiCE) routes each input token $x_{t}$ to a subset of experts, each defined in its own curvature space. First, token–expert affinities are computed in the space-like dimensions:
(15) $s_{t,j}=\mathrm{act}\!\big{(}(x_{t})_{s}^{\!\top}\,y_{j}\big{)},$
where $(x_{t})_{s}$ is the space-like projection of $x_{t}$, $y_{j}$ is the learnable centroid vector of expert $j$, and $\mathrm{act}(\cdot)$ is an activation function. Top-$K_{r}$ affinities are retained and $\ell_{1}$-normalized to give the gating weights:
(16) $g_{t,i}=\frac{\displaystyle s_{t,i}\,\mathbf{1}\!\big{[}s_{t,i}\!\in\!\mathrm{Top}_{k}(\{s_{t,j}\}_{j=1}^{N_{r}},K_{r})\big{]}}{\displaystyle\sum_{j=1}^{N_{r}}s_{t,j}\,\mathbf{1}\!\big{[}s_{t,j}\!\in\!\mathrm{Top}_{k}(\{s_{t,j}\}_{j=1}^{N_{r}},K_{r})\big{]}},$
where $N_{r}$ is the number of routed experts and $K_{r}\!<\!N_{r}$ is the routing budget. The final MiCE representation is the Lorentz-centroid combination of expert outputs:
(17) $\mathrm{MiCE}(x_{t})=x_{t}\oplus_{L}\mathrm{Centroid}\!\Bigl{(}\{\,g_{t,i}\cdot\mathrm{Expert}_{i}(x_{t},\mathcal{K}_{i})\,\}_{i\in\mathcal{A}_{t}}\Bigr{)},$
where $\mathcal{A}_{t}$ is the set of activated experts, $\mathcal{K}_{i}$ is the curvature of expert $i$, $\oplus_{L}$ denotes a Lorentz residual connection, and $\mathrm{Centroid}(\cdot)$ is the Lorentzian centroid operator. HELM further supplies hyperbolic rotary positional encodings (HOPE) and hyperbolic RMS normalization (RMSNormL), completing the toolkit for fully hyperbolic large-scale language models. This approach preserves the benefits of a fully hyperbolic network and further boosts flexibility in representation. Early results show that HELM achieves competitive or better accuracy than Hypformer on challenging reasoning tasks, with the added ability to model diverse structural patterns. The mixture-of-curvature design enables adaptive representation learning, where different experts specialize in different levels of hierarchy—from fine-grained local relationships (high curvature) to broad semantic categories (low curvature). This architectural flexibility addresses a key limitation of fixed-curvature hyperbolic models.
Furthermore, Chen et al. [16] propose fully hyperbolic frameworks based on Lorentz transformations to overcome the limitations of hybrid architectures that rely on tangent space operations. These approaches perform all operations directly in hyperbolic space, eliminating the need for exponential and logarithmic maps:
(18) $\mathbf{h}^{\prime}=\text{LorentzLinear}(\mathbf{h})=\frac{(W\mathbf{h}+b)}{\sqrt{\|W\mathbf{h}+b\|^{2}-1}}.$
This formula rescales the usual linear combination $W\mathbf{h}+b$ onto the hyperboloid (ensuring it lies in $\{x:\langle x,x\rangle_{L}=-1\}$). By operating entirely with such Lorentz-Linear layers, these models preserve hierarchy end-to-end. The Lorentz-Linear transformation maintains the hyperbolic constraint through normalization, ensuring that all intermediate representations remain on the hyperboloid manifold throughout the forward pass. This eliminates the numerical instabilities associated with repeated exponential/logarithmic mappings while preserving the geometric properties essential for hierarchical modeling. This model empirically outperforms equivalent Euclidean PLMs across a variety of NLP tasks [16].
More recently, He et al. [46] introduced HyperCore, an open-source framework providing a comprehensive suite of fully hyperbolic neural modules. These generalize standard deep learning layers (MLPs, CNNs, GNNs, and Transformers) to hyperbolic space, supporting both Lorentz and Poincaré manifolds. A key innovation is that every operation-attention, normalization, dropout, and feed-forward layers-remains within the hyperbolic manifold, avoiding tangent space projections. Specifically, HyperCore implements Lorentzian variants of core Transformer components, such as LTransformerBlock, which uses Lorentz multi-head attention and Lorentz linear layers. Each transformation $f:\mathbb{L}^{n}\rightarrow\mathbb{L}^{n}$ is designed such that outputs $f({\bf x})$ satisfy $\langle f({\bf x}),f({\bf x})\rangle_{\text{Minkowski}}=\mathcal{K}$ for curvature $\mathcal{K}<0$. With these modules, HyperCore builds the first fully hyperbolic Vision Transformer (L-ViT) and CLIP-style model (L-CLIP), validating that such architectures can effectively learn hierarchical visual semantics.
Li et al. [63] proposed HyLiFormer, which combines Poincaré hyperbolic embeddings with linear-time attention for long-sequence modeling. Motivated by skeleton-based action recognition (a hierarchical motion task), HyLiFormer projects input features into hyperbolic space using the HTC module with learnable curvature constraints. Features are unit-normalized, compressed via a hyperbolic $\tanh$ mapping, and re-embedded into the ball. The subsequent Hyperbolic Linear Attention (HLA) module uses a factorized attention mechanism based on kernelized queries and keys, computed using Lorentzian inner products and preserving the manifold structure. This architecture yields a fully hyperbolic transformer with $\mathcal{O}(N)$ attention cost and empirically outperforms Euclidean baselines [63]. The success of fully hyperbolic transformers has inspired researchers to explore alternative architectures. State-space models, with their linear scalability, present unique opportunities for hyperbolic adaptation while maintaining computational efficiency.
3.4 Hyperbolic state-space models
Beyond Transformer-based architectures, fully hyperbolic sequence models have also been explored using state-space architectures. These models address the quadratic complexity limitations of hyperbolic transformers while maintaining the ability to capture hierarchical relationships across long sequences. For instance, Hierarchical Mamba (HiM) [83], a fully hyperbolic state-space LLM that integrates the linear-time Mamba2 sequence model with hyperbolic geometry. HiM projects sequence representations into a Poincaré ball or Lorentz manifold with a learnable curvature and enforces hierarchy through specialized hyperbolic loss functions, while explicitly constraining norms to ensure numerical stability. HiM introduces explicit hierarchical constraints via learned hyperbolic geometry. It uses a learnable norm-scaling factor $\gamma$ and learnable curvature $\mathcal{K}$ so that scaling the pooled embedding by $\gamma$ before projection effectively lets the model adaptively find the optimal curvature and scale for each hierarchy level. HiM also uses a dynamic centripetal loss and clustering loss: the centripetal loss forces parent-node embeddings to lie closer to the hyperbolic origin than their children, and the clustering loss groups related points while pushing unrelated ones apart. HiM’s dynamic hyperbolic losses adapt based on learned curvature $\mathcal{K}$. The centripetal loss enforces hierarchical ordering:
(19) $L_{\text{centripetal}}=\sum_{(e,e^{+},e^{-})\in D}\max\left(\|e^{+}\|_{c}-\|e\|_{c}+\beta,0\right),$
where $e$ = anchor node embedding, $e^{+}$ = positive (parent) embedding, $e^{-}$ = negative (unrelated) embedding, $|\cdot|_{c}$ = hyperbolic norm (distance from origin in the chosen manifold), $\beta$ = curvature-dependent margin
The clustering loss groups related points while separating unrelated ones:
(20) $L_{\text{clustering}}=\sum_{(e,e^{+},e^{-})\in D}\max\left(d_{c}(e,e^{+})-d_{c}(e,e^{-})+\alpha,0\right),$
where $d_{c}(\cdot,\cdot)$ = hyperbolic distance in the chosen manifold, $\alpha$ = curvature-dependent clustering margin. The margins $\alpha$ and $\beta$ are functions of the hyperbolic radius $r=\sqrt{c}$ (where $c$ is the curvature scaling parameter), typically set as $\alpha=0.255\cdot r$ and $\beta=0.0051\cdot r$. These dynamic hyperbolic losses update their centripetal and clustering margins based on the updated radius as the curvature $\mathcal{K}$ keeps adapting. In effect, HiM’s mathematical innovation is to jointly learn a hyperbolic embedding scale and curvature while preserving tree-like structure end-to-end and scales long sequences as per state-space duality, yielding significant improvements on hierarchical reasoning benchmarks (mixed-hop prediction, multi-hop inference), compared to equivalent Euclidean models. This demonstrates that even non-Transformer architectures can be made fully hyperbolic, combining efficiency and hierarchical awareness in LLMs.
Recent advances in state-space models have led to the development of the HMamba [128], which integrates hyperbolic geometry with Mamba’s selective state-space mechanism for the sequential recommendation task. HMamba introduces two distinct architectural variants: HMamba-Full and HMamba-Half. HMamba-Full represents a fully hyperbolic architecture that processes both input and output entirely in hyperbolic space. In contrast, HMamba-Half is a partially hyperbolic variant that implements a hybrid approach using Euclidean space to balance computational efficiency with hyperbolic benefits. HMamba-Full applies a curvature-normalized discretization of the continuous-time state-space model. In practice, the discrete state matrix is computed as $\bar{A}=\exp(\Delta A\odot K(\mathcal{K}))$ with $K(\mathcal{K})=\mathrm{diag}(\sqrt{|\mathcal{K}|},1,\dots,1)$, where $\mathcal{K}<0$ is the learnable curvature parameter and $K(\mathcal{K})$ is a matrix-valued function. This new discrete state matrix $K(\mathcal{K})$ preserves the Lorentz metric under discretization. Hidden states are then evolved via Lorentz parallel transport: $h^{\prime}(i)=\text{PT}_{o\to E^{(h)}_{u,i}}(A,h(i))+BE^{(h)}_{u,i}$, where $\text{PT}_{o\to E^{(h)}_{u,i}}$ denotes the parallel transport operation from origin $o$ to point $E^{(h)}_{u,i}$ on the Lorentz manifold, ensuring that hierarchical geometry is maintained over time. These design choices make HMamba-Full fully curvature-aware. By contrast, HMamba-Half scores sequences in flat Euclidean space (using standard dot-product attention on the Mamba outputs), so it does not use any curvature-dependent operation or PTrans. This hybrid design demonstrates the trade-off between geometric expressiveness and computational efficiency: HMamba-Full achieves superior hierarchical modeling at the cost of increased computational overhead, while HMamba-Half provides a balance suitable for large-scale applications where full hyperbolic computation may be prohibitive.
Yang et al. [123] introduced SHMamba (Structured Hyperbolic State Space Model), a novel approach that combines hyperbolic geometry with state space models for audio-visual question answering (AVQA). SHMamba addresses the limitations of Euclidean space in representing hierarchical audio-visual relationships and the quadratic computational complexity of transformer-based approaches for long sequence modeling. The model leverages the Poincaré ball to project audio and visual features into hyperbolic space using an adaptive curvature hyperbolic alignment module, which dynamically adjusts curvature parameters based on concatenated modal features: $\mathcal{K}=\mathcal{K}_{0}\cdot\text{sigmoid}(\text{MLP}(K_{av}))$, where $\mathcal{K}$ is the adaptive curvature parameter and $K_{av}$ represents the concatenated audio-visual features. The structured state space component captures temporal dynamics through discrete evolution equations while maintaining global sequence information, combined with a cross fusion block that enhances cross-modal feature interaction through gating mechanisms and deep convolution operations. Experimental results demonstrate that SHMamba achieves superior performance with 78.12% fewer parameters while improving average performance by 2.53% compared to transformer-based baselines, validating the effectiveness of combining hyperbolic geometry with state space architectures for multimodal understanding tasks [123].
By operating entirely in hyperbolic space, these models better preserve hierarchical relationships throughout the network, leading to improved performance on tasks that require modeling complex hierarchical structures. However, this approach faces ongoing challenges in numerical stability and computational efficiency, particularly in maintaining geometric consistency during the discrete state evolution process. The elimination of tangent space approximations introduces new complexities in gradient computation and requires specialized optimization techniques that remain active areas of research. Having established the architectural foundations of hyperbolic LLMs, we now examine how these models have been successfully applied across diverse domains, demonstrating their practical utility and broad applicability.
4 Benchmarking and evaluation of hyperbolic LLMs
The evaluation of hyperbolic large language models presents unique challenges due to their specialized geometric properties and diverse architecture integrations of hyperbolic geometry. This section examines performance benchmarks of hyperbolic LLMs on mathematical reasoning and hierarchical language reasoning tasks, highlighting current benchmarking limitations and the need for a unified evaluation framework.
4.1 Performance benchmarks across HypLLMs
(1) Mathematical reasoning and arithmetic tasks
Hyperbolic fine-tuned models demonstrate exceptional performance on complex reasoning tasks. Table 3 presents evaluation results of HypLoRA and baseline models across multiple mathematical reasoning datasets. The results demonstrate that HypLoRA consistently outperforms both standard LoRA and DoRA across complex reasoning tasks, with particularly significant improvements on the most challenging AQuA dataset (13.0% improvement for Gemma-7B), validating the effectiveness of hyperbolic geometry for mathematical reasoning. Table 4 presents hyperbolicity measurements across different model scales and datasets, revealing consistent patterns that support hyperbolic geometric assumptions.
Table 3: Mathematical reasoning performance: HypLoRA vs. baseline methods (Accuracy %)  [119].
Backbone Model MAWPS [56] SVAMP [82] GSM8K [25] AQuA [66]
LLaMA-7B LoRA 79.0 52.1 37.5 18.9
DoRA 80.0 48.8 39.0 16.4
HypLoRA 79.0 49.1 39.1 20.5
LLaMA-13B LoRA 83.6 54.6 47.5 18.5
DoRA 83.0 54.6 18.9
HypLoRA 83.2 54.8 49.0 21.5
Gemma-7B LoRA 91.6 76.2 66.3 28.9
DoRA 91.7 75.9 65.4 27.7
HypLoRA 91.5 78.7 69.5 32.7
LLaMA3-8B LoRA 92.7 78.9 70.8 30.4
DoRA 92.4 79.3 71.3 33.1
HypLoRA 91.6 80.5 74.0 34.2
Table 4: $\delta$-Hyperbolicity analysis across LLM architectures and datasets [119].
Backbone MAWPS [56] SVAMP [82] GSM8K [25] AQuA [66]
LLaMA-7B $08\pm 0.02 0.0$$09\pm 0.01$ $\pm 0.01 0.10±0.01$$\pm 0.01$
LLaMA-13B $08\pm 0.01 0.0$$9\pm 0.01 0.0$$09\pm 0.01$ $0.10\pm 0.01$
Gemma-7B $1\pm 0.01 0.11±0$$1\pm 0.01$ $0.11\pm 0.01$ $0.12\pm 0.01$
LLaMA3-8B $06\pm 0.01 0.0$$7\pm 0.01 0.0$$7\pm 0.01 0.0$$08\pm 0.01$
Average $08\pm 0.01 0.0$$9\pm 0.01 0.0$$09\pm 0.01$ $0.10\pm 0.01$
(2) Hierarchical language reasoning (mixed-hop prediction) task
Hierarchical language reasoning tasks evaluate models’ ability to understand and predict relationships across different levels of semantic hierarchies. The mixed-hop prediction task is particularly challenging as it requires models to predict the exact hierarchical distance between entities in ontological structures, distinguishing between direct (1-hop) and indirect multi-hop relationships. Hierarchical Mamba (HiM) [83] demonstrates exceptional performance in this domain by integrating efficient Mamba2 state-space models with hyperbolic geometry. Table 5 presents our experimental evaluation results for mixed-hop prediction tasks across ontological datasets, comparing hyperbolic models against the Euclidean baseline SentenceMamba-16M [83]. SentenceMamba-16M is a Mamba2-based large language model with 16 million parameters designed to generate high-quality sentence embeddings. Notably, HiM-Poincaré achieves the best performance across all datasets, highlighting the effectiveness of hyperbolic geometry combined with state space models for capturing hierarchical language structures. It should be noted that the original HiT uses a pretrained model (e.g., all-MiniLM-L6-v2) for hyperbolic space training with curvature $\mathcal{K}=-1/384$, which is slightly Euclidean-like. In our experiments, we use HiT-Rand-Init with random initialization and higher curvature $\mathcal{K}=-1.0$ for more hyperbolic behavior. We use this randomly initialized variant of HiT for fair comparison with HiM, since HiM is also trained with random initialization.
Table 5: Mixed-hop prediction performance (F1 score %): Comparison of Hierarchical Mamba (HiM) and Hierarchical Transformer (HiT) with Euclidean baseline SentenceMamba-16M across ontological datasets.
Backbone Model WordNet DOID
SentenceMamba-16M (Mamba-2 based) Euclidean baseline 61.5 43.6
all-MiniLM-L6-v2 (Transformer-based) HiT-Rand-Init Poincaré 84.6 83.7
SentenceMamba-16M (Mamba-2 based) HiM-Poincaré 85.9 90.2
4.2 Benchmark limitations and future directions
Existing benchmarking approaches reveal several limitations that need to be addressed in future research:
(1) Hierarchical structure evaluation: Existing benchmarks primarily focus on task-specific accuracy rather than measuring the quality of hierarchical representations. Future benchmarks should incorporate metrics that explicitly evaluate hierarchical structure preservation and multi-scale reasoning capabilities.
(2) Scalability benchmarks: Limited evaluation of hyperbolic models on large-scale datasets restricts understanding of scalability characteristics. Future benchmarking should include systematic evaluation across varying dataset sizes and model scales.
(3) Unified evaluation framework: The absence of standardized evaluation protocols across HypLLM categories complicates comparative analysis. Development of unified benchmarking frameworks would enable more robust cross-model comparisons and accelerate research progress.
The empirical evidence consistently demonstrates that hyperbolic geometry provides significant advantages for tasks involving hierarchical structures, complex reasoning, and multi-modal understanding, while revealing important areas for continued research and development in benchmarking methodologies.
5 Emerging applications of HypLLMs
The theoretical foundations and architectural innovations in hyperbolic geometry have enabled HypLLMs to demonstrate remarkable versatility across diverse application domains. As illustrated in Figure 6, hyperbolic large language models have found successful applications spanning computer vision, natural language processing, multimodal representation learning, neuroscience, and biomedical domains. Each application domain leverages different aspects of hyperbolic geometry’s unique properties—from capturing visual concept hierarchies in computer vision to modeling temporal branching patterns in sequence data, and from representing cross-modal semantic relationships to uncovering latent organizational principles in neural networks. This section examines how HypLLMs have been successfully deployed across various domains, demonstrating both the practical utility and broad applicability of hyperbolic geometric principles in LLM architectures.
Figure 6: Applications of hyperbolic large language models (HypLLMs) across different domains. Each bubble corresponds to a domain, with included HypLLMs and their corresponding authors.
5.1 HypLLMs for computer vision
Large-scale vision tasks often involve hierarchical relationships, for example, object recognition can be organized by a taxonomy of categories (from general classes to fine-grained subtypes). Hyperbolic representations have begun to show promise in computer vision by capturing these inherent hierarchies. One prominent approach is to incorporate hyperbolic geometry into vision-language models. Srivastava et al. [94] proposed PoinCLIP, a vision–language model that projects CLIP embeddings into the Poincaré ball to better reflect the conceptual hierarchy between images and text descriptions. By learning joint image-text representations entirely in a hyperbolic space, PoinCLIP preserves the natural “is-a” relationships (e.g., an image of a dog is also an image of an animal) as distances in the embedding space. Using hyperbolic image-text embeddings yields improved zero-shot classification and image retrieval performance over the standard (Euclidean) CLIP, especially for abstract or coarse-grained (hierarchical) categories. Building upon these foundational insights, more comprehensive hyperbolic vision-language architectures have emerged. Mandica et al. [72] extend the BLIP-2 architecture into hyperbolic space to capture hierarchical image-text relationships called Hyperbolic BLIP-2. Their model uses a frozen image encoder and frozen LLM but projects visual and language embeddings into the Poincaré ball using the exponential map. Contrastive learning is performed with cosine similarity (instead of geodesic distance) to avoid instability. Two regularization methods-Random Query Selection (RQS) and Random Text Pruning (RTP)-encourage diversity and robustness in hyperbolic embedding norms. This results in meaningful norm-based uncertainty estimates and strong retrieval/captioning performance.
Qu et al. [87] propose LLM-AR, a framework that projects skeleton sequences into a linguistic format for LLM-based action recognition, utilizing a VQ-VAE with a hyperbolic codebook for the encoder. While the LLM component in LLM-AR operates in Euclidean space, the use of hyperbolic geometry in the codebook highlights the growing convergence of hyperbolic representation learning and LLMs for structured vision tasks. In parallel, HyLiFormer [63] demonstrates the advantages of fully learning in hyperbolic space for skeleton-based action recognition. Beyond vision-language alignment, hyperbolic embeddings have also been applied to purely visual tasks. For instance, image classifiers can be enhanced by replacing the final Euclidean feature layer with a hyperbolic space that encodes class hierarchies. This allows the model to naturally leverage hierarchical label structures, e.g., an image misclassified within the correct superclass is penalized less, leading to better generalization on imbalanced or hierarchical datasets [30].
5.2 HypLLMs for sequence modeling
Many sequential datasets have an intrinsic hierarchical organization, where hyperbolic geometry can be leveraged to model long-range dependencies or branching patterns. Hyperbolic representations can model branching temporal or logical sequences more faithfully than Euclidean ones. Klimovskaia et al. [54] introduced Poincaré maps to analyze single-cell gene expression sequences. In single-cell developmental trajectories, cells differentiate following a tree-like hierarchy with multiple branches. They showed that mapping high-dimensional scRNA-seq data into a 2D hyperbolic disk uncovers these branching hierarchies more faithfully than Euclidean PCA or UMAP, while preserving meaningful distances between cell states. Each Poincaré map (hyperbolic embedding) effectively acts as a continuous analog of a lineage tree, enabling clearer visualization and more accurate clustering, pseudotime ordering, and lineage inference. This demonstrates the power of hyperbolic representations for sequence modeling in biology, where chronological or evolutionary sequences form implicit hierarchies. Beyond biology, hyperbolic sequence modeling has potential in natural language processing: for instance, modeling discourse or narrative structure as a hierarchy, or encoding the latent tree structure of sentences in a sequence model’s state space.
Recent work has begun applying hyperbolic LLM techniques to sequential data in NLP and beyond, using novel and efficient architectures. The HiM model [83] exemplifies how hyperbolic geometry can enhance language sequence modeling. By leveraging Mamba’s selective state-space mechanism, HiM projects outputs into a curved space (Poincaré or Lorentz) with learnable curvature and dynamic margins. Across various ontological datasets, HiM captures multi-level linguistic hierarchies more effectively than Euclidean baselines. Similarly, in the recommendation domain, Zhang et al. [128] introduce Hyperbolic Mamba (HMamba) to model user-item sequential interactions. HMamba unifies Mamba’s linear-time state-space architecture with hyperbolic representational power, implementing a curvature-aware state space and stabilized Riemannian operations to maintain hierarchy in user preference sequences. Thanks to this design, HMamba achieves 3–11% higher top-$K$ recommendation accuracy than comparable Transformer-based methods, while retaining Mamba’s $\mathcal{O}(L)$ efficiency for long interaction streams. Ma et al. [70] proposed HERec, a hyperbolic graph-LLM framework that unifies semantic information from LLMs with collaborative user-item graphs in hyperbolic space. HERec introduces a hierarchical representation structure for user-adjustable exploration-exploitation trade-offs, achieving state-of-the-art performance on both utility and diversity metrics in recommendation tasks. These cases illustrate how HypLLM approaches can be extended to various sequence learning problems-from natural language to recommender systems-capturing latent hierarchical structures without sacrificing scalability.
Furthermore, the benefits of hyperbolic state-space models likely extend to other sequential domains. Pandey et al. [80] report that for dynamic graph sequences (i.e., temporal networks), Mamba-based models can match or exceed Transformer performance in link prediction tasks while being significantly more efficient. Although their work does not incorporate hyperbolic geometry, it highlights the potential of state-space architectures in temporal graph modeling, suggesting that future work could explore integrating hyperbolic embeddings into such models to better capture evolving hierarchical structures in dynamic networks. In summary, HypLLM approaches for sequence data exploit hyperbolic space’s ability to naturally embed tree-like progression, whether it is cell differentiation lineages or abstract syntax trees in text, leading to richer and more interpretable sequence representations.
5.3 HypLLMs for multimodal representation learning
Hyperbolic LLM techniques are also being applied to multimodal representation learning problems, where data from different modalities (text, vision, graphs, etc.) have underlying hierarchical correspondences. Xiong et al. [116] propose a model called HyperSurv that fuses pathology images and text reports for cancer survival prediction. In this task, doctors’ free-text reports describe visual findings in whole-slide images (WSIs); the descriptions have a hierarchical structure (e.g., broad terms like “tumor” entail specific attributes like “necrosis” or “mitosis count”). HyperSurv maps both the image features and the report embeddings into a shared hyperbolic space, using hyperbolic cones to enforce entailment constraints (so that a general concept in text is located “above” the more specific visual features it refers to). This approach naturally captures the one-to-many relationships between high-level medical terms and multiple image regions. As a result, the model achieved state-of-the-art accuracy in survival outcome prediction across several cancer datasets, outperforming Euclidean baselines by effectively modeling the multi-modal hierarchy of concepts and images. Building on similar principles for cancer survival prediction, Yang et al. [117] introduced HySurvPred, which integrates histopathology images and genomic data in hyperbolic space for survival prediction. Unlike HyperSurv’s focus on pathology images and text reports, HySurvPred specifically addresses the hierarchical structures in both histopathology data (from tissue to cell scales) and genomics data (from biological networks to genes). The framework introduces three key innovations: Multimodal Hyperbolic Mapping (MHM) to capture intra-modal hierarchies, Angle-aware Ranking-based Contrastive Loss (ARCL) to preserve the ordinal nature of survival time, and Censor-Conditioned Uncertainty Constraint (CUC) to effectively utilize censored data. The framework operates by mapping multimodal features from Euclidean space into hyperbolic space to better capture hierarchical relationships, achieving superior performance on five TCGA benchmark datasets and demonstrating the effectiveness of hyperbolic multimodal fusion for complex medical prediction tasks [117].
Another emerging direction is e-commerce and web multimodal data: Amazon researchers have explored hyperbolic embeddings for product catalogs, which include text descriptions, images, and user interaction graphs. By learning visuo-lingual concept embeddings in a hyperbolic space, a system can represent, say, a product image and its title within a shared hierarchy (where general product categories branch into specific products). Choudhary et al. [21] proposed the ANTHEM (Attentive Hyperbolic Entity Model) framework, which models queries and products as hyperboloids in hyperbolic space and uses attention mechanisms to capture hierarchical and compositional relationships. Evaluated on large-scale e-commerce data, ANTHEM demonstrated more than 10% improvement over state-of-the-art product search methods, highlighting the benefits of hyperbolic geometry for representing product knowledge graphs and “also-viewed” relationships with less distortion and in fewer dimensions. Cheng et al. [18] introduced HyperLLM, a model-agnostic framework that integrates large language models with hyperbolic space to capture hierarchical information in recommender systems. HyperLLM leverages LLMs to generate multi-level classification tags and aligns user-item interactions in hyperbolic space through contrastive learning, achieving over 40% improvement compared to prior methods. Cheng et al. [18] further validate HyperLLM’s effectiveness through detailed comparisons against multiple strong baselines. This work demonstrates the synergy between LLMs and hyperbolic geometry for hierarchical recommendation and training stability.
Similarly, the PoinCLIP model discussed in Section 5.1 integrates hyperbolic geometry to better align images with text by their conceptual hierarchy, improving zero-shot recognition of image–text pairs [94]. Hyperbolic BLIP-2 discussed earlier is one of the first billion-parameter hyperbolic multimodal LLMs [72]. In the audio-visual domain, Yang et al. [123] proposed SHMamba for audio-visual question answering, demonstrating how hyperbolic geometry can enhance multimodal understanding in sequential data. The model addresses the inherent hierarchical structure present in audio-visual content, where sounds and visual elements often follow tree-like organizational patterns (e.g., musical instrument categories branching into specific instruments, or object taxonomies in visual scenes). SHMamba’s hyperbolic alignment loss $L_{\text{align}}=\|\mathbf{W}_{v,\text{norm}}-\mathbf{W}_{a,\text{norm}}\|^{2}$ enforces semantic consistency between audio and visual modalities in the Poincaré ball, where $\mathbf{W}_{v,\text{norm}}$ and $\mathbf{W}_{a,\text{norm}}$ represent normalized similarity matrices computed in the hyperbolic tangent space. The adaptive curvature mechanism allows the model to flexibly explore different hierarchical structures depending on the specific audio-visual content characteristics, making it particularly effective for complex multimodal scenes with varying levels of semantic hierarchy. Evaluated on MUSIC-AVQA and AVQA datasets, SHMamba achieved state-of-the-art performance while maintaining computational efficiency, highlighting the potential of hyperbolic state space models for real-time multimodal applications. These examples illustrate that for multimodal problems involving inherently hierarchical data (whether medical ontologies linking text and images, or product ontologies linking descriptions, images, and user behavior), hyperbolic LLMs provide a powerful tool. They enable a unified embedding space where different modalities meet, and where distance directly encodes semantic or conceptual similarity across modalities, something especially beneficial when the modality interactions are complex and multi-scale.
5.4 Hyperbolic embeddings for brain network analysis
Beyond language and knowledge graphs, hyperbolic embeddings have recently been applied in neuroscience to model the brain’s network organization. The concept of neural manifolds underlying brain function provides a crucial theoretical foundation for understanding why hyperbolic embeddings are particularly effective for neural data. Gallego et al. [37] demonstrated that neural population activity during movement is constrained to low-dimensional manifolds spanned by specific patterns of correlated neural activity, termed “neural modes.” This manifold-based view of neural computation suggests that the coordinated activity of neural populations can be captured by relatively few covariation patterns, which naturally aligns with the hierarchical representational capacity of hyperbolic space. Furthermore, Fortunato et al. [36] provided compelling evidence that these neural manifolds are intrinsically nonlinear, particularly during complex behaviors that require varied neural activity patterns. Their analysis across monkey, mouse, and human motor cortex revealed that nonlinear manifolds consistently outperform flat manifolds in capturing neural population dynamics, with the degree of nonlinearity varying across architecturally distinct brain regions. This finding has profound implications for hyperbolic neural models, as it suggests that the curved geometry of hyperbolic space may be fundamentally better suited to represent the intrinsic nonlinear structure of neural population activity than traditional Euclidean approaches.
Baker et al. [5] presented a framework for embedding MEG (magnetoencephalography) functional brain networks into hyperbolic space. The motivation is that brain connectivity graphs possess a hierarchical structure (from local circuits to distributed modules) which might be captured more naturally in a hyperbolic space $\mathbb{H}^{n}$. Indeed, in their study, high-dimensional brain networks were mapped into a low-dimensional hyperbolic space while preserving both local and global geometric relationships with lower distortion than competing Euclidean embedding methods. This means that even in 2 or 3 dimensions, the hyperbolic representation retained important network properties (connectivity patterns, community structure) that Euclidean PCA or spectral embeddings struggled to maintain. Baker et al. define a new metric in the latent space-the radial coordinate of a node’s embedding, as a proxy for that node’s hierarchical importance in the brain network [5]. Using this metric, they discovered subtle but consistent differences between healthy individuals and those with subjective cognitive decline (SCD). In particular, participants with early cognitive decline showed an increased hierarchy (greater average embedding radius) in specific subnetworks (e.g., dorsal and ventral attention networks) relative to healthy controls. These results suggest that the brains of SCD subjects exhibit a shift toward a more treelike, hierarchical organization, which hyperbolic embeddings can detect.
In addition to cognitive decline detection, hyperbolic neural models have recently been employed to model aging trajectories in large-scale brain connectivity data. Ramirez et al. introduced a Fully Hyperbolic Neural Network (FHNN) that operates directly in Lorentz space to embed MEG-derived brain graphs from over 500 subjects [90]. The model captures subtle, hierarchical reorganizations in subnetworks (e.g., attention, default mode networks) that occur with aging. This framework leverages the radius of Lorentz embeddings to quantify hierarchy and reveals that elderly individuals show a reduction in embedding hierarchy-opposite to the increase observed in early cognitive decline.
This application underscores the versatility of hyperbolic representations: they are not only mathematically appealing for theoretical data hierarchies, but also empirically powerful in analyzing real biological networks. By capturing multi-scale structure in brain connectivity data, hyperbolic LLM techniques can potentially contribute to cognitive modeling and clinical neuroscience. Such cross-disciplinary success strengthens the case that hyperbolic geometry is a unifying language for complex hierarchical systems across domains.
6 Discussion
The evolution of hyperbolic geometry from mathematical abstraction to practical computational tool represents a paradigm shift in representation learning for hierarchical data. Our taxonomy reveals distinct trade-offs across the four HypLLM categories: (1) hybrid hyperbolic-Euclidean models strategically integrate curved geometry into existing architectures through exponential/logarithmic mappings, proving effective for knowledge graph completion and taxonomy reasoning; (2) hyperbolic fine-tuned models provide parameter-efficient adaptation of pre-trained models, particularly successful in boosting logical reasoning and arithmetic problem-solving; (3) fully hyperbolic models operate entirely in curved space with specialized geometric operations, aimed at improving general NLP benchmark performance while maintaining computational efficiency; (4) hyperbolic state-space models combine the linear complexity of state-space architectures with hierarchical modeling capabilities, enabling efficient processing of long sequences with complex hierarchical structures.
Beyond traditional NLP and computer vision applications, recent neuroscience research has demonstrated that hyperbolic geometry provides fundamental insights into human cognitive processes and perception. Recent advances in Bayesian hyperbolic embedding methods [86] have enabled more principled approaches to modeling complex hierarchical data structures found in neural systems. A particularly promising direction lies in leveraging insights from low-dimensional, intrinsically nonlinear manifolds that provide more natural representation space for neural data [37, 36]. Lee et al. [60] revealed that object concept memorability follows hierarchical structures better captured by hyperbolic rather than Euclidean geometry, showing that memorable objects cluster near the center of hyperbolic representational spaces while forgettable objects scatter toward the periphery. This finding suggests that human memory systems may inherently operate within curved geometric frameworks that preserve hierarchical relationships among concepts. Similarly, Zhou et al. [131] demonstrated that human olfactory perception can be accurately modeled using three-dimensional hyperbolic spaces, where natural odors and perceptual descriptions both exhibit hierarchical organization that mirrors the biochemical networks producing these compounds. The cross-disciplinary validation of HypLLMs strengthens the theoretical foundation of hyperbolic geometry as a unifying framework for complex hierarchical systems. The remarkable success across diverse domains—detecting subtle patterns in brain networks, improving cancer outcome prediction, and enhancing multiscale reasoning in NLP—validates its ability to be a versatile tool across AI fields. This broad applicability suggests that hyperbolic principles may represent universal organizational patterns in both biological and artificial intelligence systems.
7 Conclusion
Hyperbolic geometry offers a compelling framework for advancing large language models, enabling them to capture hierarchical structures that are elusive in Euclidean space. This paper has provided a thorough discussion of hyperbolic large language models (HypLLMs) from modest hyperbolic embedding tweaks to fully Lorentzian Transformers, exploring their theoretical foundations, architectural approaches, and emerging applications. As demonstrated through our architectural taxonomy (Table 1), and cross-domain applications (Figure 6), hyperbolic geometry provides a unifying mathematical framework that spans from hybrid integration approaches to fully curved-space architectures, with successful deployment across diverse fields including natural language processing, computer vision, multimodal representation learning, neuroscience, and biomedicine. We highlight the significant potential of hyperbolic geometry for enhancing the representation capabilities of large language models, particularly for data with inherent hierarchical structures. Crucially, hyperbolic approaches have demonstrated broad utility, powering improvements in natural language understanding, vision–language grounding, biomedical knowledge graphs, and even neuroscience.
Despite these advances, several open challenges remain. Continued research on stabilization, such as multi-precision floating points and curvature-sensitive initialization, will be critical as models scale up. Looking ahead, we anticipate several transformative directions for HypLLMs. Hybrid-curvature architectures, such as mixture-of-curvature experts in models like HELM, and learnable-curvature architectures such as HiM, may unlock new capabilities by representing different aspects of data with optimal geometric spaces. Numerical stability for deep hyperbolic models, efficient large-scale training, and the development of unified benchmarks for hierarchical and multi-scale reasoning represent critical directions for future research.